OpenStudy (waheguru):

How Could You Factor This 2(a+b)^2+5(a+b)+3

OpenStudy (mathstudent55):

Replace (a + b) with x. Does this look familiar now?

OpenStudy (waheguru):

Hmm, Ill try that

OpenStudy (waheguru):

2x^2+5x+3

OpenStudy (waheguru):

@mathstudent55 What can I do from here

OpenStudy (mathstudent55):

Do you know how to factor that? You need two sets of parentheses: ( )( )

OpenStudy (waheguru):

Yea but there is 5x and what combination adds to 5 that also has a product of 6?

OpenStudy (mathstudent55):

The numbers here are small so we can use the guess method. The 2x^2 term is 2x * x, right? So it goes here: (2x )(x )

OpenStudy (waheguru):

Is there a mathematical way rather than guessing?

OpenStudy (mathstudent55):

The term 3 is 3 * 1, so now you need to look at only two possibilities: (2x + 3)(x + 1) (2x + 1)(x + 3)

OpenStudy (mathstudent55):

Yes, I can show you the mathematical way also.

OpenStudy (waheguru):

Yea, I would like to see the mathematical way

OpenStudy (waheguru):

How can we factor 2x^2+5x+3

OpenStudy (mathstudent55):

Let's start from the beginning. We are factoring a trinomial of the form \(ax^2 + bx + c\) Your trinomial is \(2x^2 + 5x + 3\) You have a = 2, b = 5, and c = 3, ok?

OpenStudy (mathstudent55):

1. Multiply ac together. For you, ac = 2*3 = 6.

OpenStudy (waheguru):

yea, just got that

OpenStudy (mathstudent55):

2. Find two factors of ac whose sum is b. In your case, you need two factors of 6 whose sum is 5.

OpenStudy (waheguru):

Yep and the final answer which I got was (2a+2b+3)(a+b+1)

OpenStudy (mathstudent55):

3. Break up the middle term with the two factors you found. In your case, the factors are 2 and 3. Now you write 2x^2 + 2x + 3x + 3 4. Factor by parts. That is, factor a common factor out of the first two terms amd a common factor out of the last two terms. In your case, 2x(x + 1) + 3(x + 1) 5. Factor out the common term. In your case: (x + 1)(2x + 3)

OpenStudy (mathstudent55):

Now we substitute back the a + b for x that we did earlier on. (x + 1)(2x + 3) = (a + b + 1)[2(a + b) + 3] = (a + b + 1)(2a + 2b +3) Which is what you got. Good job!